if a model rocket was measured by the vertical motion model h=-16t2+96t, how many seconds would it take for the model rocket to reach 80 feet for the second time?
My math teacher asked a similar vertical motion question and if i see this one worked out i might understand it better.

Answer by benn143(2)

First set it up like it says.
80=-16t^2+96t
then subtract the 80 from both sides.
0=-16t^2+96t-80
then divide the whole thing by -16.
0=t^2-6t+5
then you simply factor and solve
0=(t-5)(t-1)
t-5=0 t-1=0
t=5 or t=1
Now these answers are when the rocket hit 80 feet on its way up and way down.
So the answer would be 5 seconds because it takes longer to come down.
Hope that helps!